frequen temper what could be the length of pipe B? A, which is 3.50 m long and open at both ends, oscillates at its second harmonic uency mode. Pipe B, which is closed at one end, oscillates at its third harmonic ency mode. This frequency of B happens to match the frequency of A. If the air erature is 25°C A. What is the fundamental frequency of pipe B? B.
Answer #1

Its given that

fA = fB

nA v/2 La = nB v/4 Lb

Lb = nB La/2nA

nA = 2 ; nB = 3 ; La = 3.5

Lb = 3 x 3.5/(2 x 2) = 2.625 m

Hence, Lb = 2.625 m

B)f = v/4L

v at 25 deg C = 346.75 m/s

f = 346.75/(4 x 2.625) = 66.05 Hz

Hence, f = 66.05 Hz